OpenStudy (anonymous):

six times a certain number minus eight has the same value as four times the number plus two. what is the number?

Directrix (directrix):

Let x be the certain number. 6x - 8 = 4x + 2. 2x = 10 x = 5

OpenStudy (anonymous):

what about 3b-13+4b=7b+1

Directrix (directrix):

What about it? :)

Directrix (directrix):

3b-13+4b=7b+1 7b -13 = 7b + 1 There can be no solution. There is no answer so do NOT write 0 as the answer.